OpenStudy (anonymous):

Solve. 6x^2 + x = 12

OpenStudy (anonymous):

(3x-4)(2x+3)=0 x=4/3 or -3/2

OpenStudy (anonymous):

6x^2+x-12=0 6x^2+9x-8x-12=0 3x(2x+3)-4(2x+3)=0 (3x-4)(2x+3)=0

OpenStudy (anonymous):

I UNDERSTAND NOW. THANK YOU.